OpenStudy (anonymous):

multiply and simplify y^3 * y^-6 * y^-2 * y A. y^-4 B. 1/y^4 C. y^-5 D. 1/y^3 please help?

OpenStudy (luigi0210):

\[y^3*y^{-6}*y^{-2}*y\]

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

\[y^{-1}=\frac{ 1 }{ y }\]

OpenStudy (anonymous):

ok?

OpenStudy (luigi0210):

Use that to change everything.

OpenStudy (anonymous):

idk how to do these

OpenStudy (luigi0210):

\[y^3*\frac{ 1 }{ y^6 }*\frac{ 1 }{ y^2 }*y\]

OpenStudy (anonymous):

im lost..

OpenStudy (luigi0210):

On what part?

OpenStudy (anonymous):

on all of it

OpenStudy (luigi0210):

When you have a negative exponent you have to flip it so it can have a positive exponent

OpenStudy (anonymous):

is it a? did the math

OpenStudy (luigi0210):

Yes it is, good job

OpenStudy (anonymous):

B